OpenStudy (anonymous):

Can someone please help step by step? d/dx [ ln(1- e^2x) / e^x^2]

zepdrix (zepdrix):

Do you remember this one?\[\Large\bf \frac{d}{dx}\ln x\quad=\quad ?\]

OpenStudy (anonymous):

I do not. But is that just x?

zepdrix (zepdrix):

Oh boy.. you have a bunch of previous questions involving the natural log. I would hope you have that one memorized by now :( Ok so this is an important one to remember:\[\Large\bf \frac{d}{dx}\ln x\quad=\quad \frac{1}{x}\]That will help us when we need to take the derivative of the numerator.

zepdrix (zepdrix):

\[\large\bf \left(\frac{\ln(1- e^{2x})}{e^{x^2}}\right)'\quad=\quad \frac{\color{#DD4747}{(\ln(1- e^{2x}))'}e^{x^2}-\ln(1-e^{2x})\color{#DD4747}{(e^{x^2})'}}{(e^{x^2})^2}\]

zepdrix (zepdrix):

So we'll need to start with the `Quotient Rule`. There is our setup.

OpenStudy (anonymous):

ok. Im with you so far

zepdrix (zepdrix):

So let's see if we can take the derivative of the first red part.

zepdrix (zepdrix):

\[\Large \left(\ln[\color{#DD4747 }{stuff}]\right)'\quad=\quad \frac{1}{\color{#DD4747 }{stuff}}(\color{#DD4747 }{stuff})'\]So after we apply the rule for differentiating natural log, we'll have to apply the chain rule, multiplying by the dervative of the inner function.

zepdrix (zepdrix):

\[\Large \left(\ln[\color{#008353 }{1-e^{2x}}]\right)'\quad=\quad \frac{1}{\color{#008353 }{1-e^{2x}}}(\color{#008353 }{1-e^{2x}})'\]

OpenStudy (anonymous):

ohh ok.

OpenStudy (anonymous):

I thought ln 1 was zero/

zepdrix (zepdrix):

Mmm well we have a bunch of other stuff in the log with the 1, so we can't just pull apart the ln1 from it :\

OpenStudy (anonymous):

ohh that makes sense.

OpenStudy (anonymous):

with you now

zepdrix (zepdrix):

So applying the chain rule, we need the derivative of that thing with the prime on it.\[\Large (1-e^{2x})'\quad=\quad (0-e^{2x}(2x)')\]

zepdrix (zepdrix):

We have yet another chain to be applied D: The exponential gives us the same thing back, but with the chain rule applied.

OpenStudy (anonymous):

oh ok. Sorry for delay

OpenStudy (anonymous):

What is the next step?